I am trying to help my daughter with homework. Since it has been many years I have forgotten how to do. How to i anwer this ...Seperate $89 into two parts so that the second part is $4less than twice the first part.
Answer by checkley71(8403)

89=X+(2X-4) OR 89=X+2X-4 OR 89+4=3X OR 93=3X OR X=93/3 OR X=31 & 2X-4=62-4=58
PROOF 31+58=89 OR 89=89 SO THE TWO NUMBERS THAT MEET THESE REQUIREMENTS ARE 31 & 58